The equipment is expected to work 90,000 hours during its 5 year useful life with a salvage value of $4,700. The productivity level is as follows:

Year 1 - 30,000 hours; Year 2 - 25,000 hours, Year 3 - 15,000 hours, Year 4 - 12,000 hours, Year 5 - 8,000 hours
Record the depreciate expense for year 3
Prepare the partial balance sheet at the end of year 3
The truck was disposed of on January 1, Year 5. Prepare the journal entry to record the disposal of the truck
Equipment/Machine
Purchase Price          45,000
Sales Taxes           3,900
Freight Insurance              600
Shipping/transport costs           1,700
Installation and Testing           1,500
Training costs           2,000
Total Acquisition cost           54,700
Answer #1

As per Ind AS 16 an item of property, plant and equipment should initially be recorded at cost. Cost includes all costs necessary to bring the asset to working condition for its intended use. This would include not only its original purchase price but also costs of site preparation, delivery and handling, installation, related professional fees for architects and engineers, and the estimated cost of dismantling and removing the asset and restoring the site.

Costs that do not provide future economic benefits are expensed in the period incurred. The initial training costs are not necessary to get the asset ready for use. Rather, the training costs are necessary to get the employees ready to use the asset. Thus, the training costs are immediately expensed.

Total Acquisition Cost will be:

Purchase Price 45000
Sales Taxes 3900
Freight Insurance 600
Shipping/transport costs 1700
Installation and Testing 1500
Total Acquisition cost 52700

Depreciable Value = Acquisition cost - Salvage value at end

= 52700 - 4700

= 48000

Total Hours worked by Machine in its life time = 90000 hours

Depreciation expenses for Year 3 :

Year 3 prodution unit = 15000 Hours

Depreciation will be = Depreciable amount * Hours worked / Total hours

= 48000*15000/90000

= $ 80000

Balance Sheet at end of Year 3

Machine A/c = 52700

Accumulated Depreciation = 37,333.33

Journal entry for disposal of Asset:

Accumulated Depreciation A/c

48000

Cash/Bank/Party A/c

4700

To Asset A/c 52700
(Being Asset sold and Accumulated depreciation a/c adjusted)

Hope the above meet your understanding.

Do let me know in case of any clarification.

Give it a thumbs up as it is the best compliment to me.